A certain recipe calls for 2/5 cup of flour. A chef is making an amount that is 5/7 of the recipe. How much flour should the chef use?
Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

5/7 * 2/5 = 10/35 = 2/7 of a cup
of flour.
Hope this helps.
:-)

Answer by stanbon(75887)

A certain recipe calls for 2/5 cup of flour. A chef is making an amount that is 5/7 of the recipe. How much flour should the chef use?
----
Ans: (5/7)(2/5) = 2/7 cup of flour
===================================
Cheers,
Stan H.
============